OpenStudy (anonymous):

Simplify this radical and remember that i =sqrt (-1) problem: √(-17) *√(-48)

OpenStudy (anonymous):

Hint:\[\sqrt{-17}\sqrt{-48}=i\sqrt{17}\ i\sqrt{48}=\cdots\]